OpenStudy (anonymous):

the qutionet of a number and -7 decreased by 2 is 10 find the number

OpenStudy (anonymous):

just answered this in another post. did you get it?

OpenStudy (anonymous):

this is another one

OpenStudy (anonymous):

no i just did it but i will be happy to do it again

OpenStudy (anonymous):

\[\frac{x}{-7}-2=10\] \[\frac{x}{-7}=10+2\] \[\frac{x}{-7}=12\] \[x=-7\times 12\] \[x=-84\]

OpenStudy (anonymous):

-100/-50